Solve the inequality.
((3x)/(7-x)) > x

Answer by RAY100(1637)

(3x) /(7-x) > x
3x > x(7-x)
3x> 7x -x^2
0> -x^2 +4x
0> x(-x +4)
.
x<0,,,,x>4
.
check ,,x<0,,use x=-1, -3/8 > -1,,,,,ok
.
0less than x less than 4,,use x= 1,,,3/6>1,,,,,,not true
.
x>4,,,use x=5,,,15/2 > 5,,,,,ok
.
therefore ,,,x<0,,,x>4
.